Ching leaves San Francisco and travels toward Los Angeles at 50 mi/h. An hour later, Phil leaves Los Angeles and travels toward San Francisco at 60 mi/h. If the two cities are 380 miles apart, how many hours will it take for Ching to meet Phil?
Answer by ptaylor(2198)

Distance(d)equals rate(r) times time(t) or d=rt; r=d/t and t=d/r
Let t=number of hours its takes for Ching to meet Phil after Phil leaves
Then t+1=total number of hours it will take for Ching to meet Phil
distance Ching travels=50t+50
distance Phil travels=60t
Now when the sum of the distances traveled by each equals 360 miles, then they will be together. So
50t+50+60t=360 subtract 50 from both sides
50t+50-50+60t=360-50 collect like terms
110t=310 divide both sides by 110
t=2.818 hours---hours it takes for Ching to meet Phil after Phil leaves
t+1=2.818+1=3.818 hours --total hours it takes for Ching to meet Phil
Check
60*2.818+50*3.818=360
169.08+190.9=360
~360=360

Hope this helps---ptaylor